OpenStudy (anonymous):

how do i solve this? 7^2t-1=35

OpenStudy (helder_edwin):

do you mean \[ \Large 7^{2t}-1=35 \]

OpenStudy (anonymous):

no it's 2t-l in the exponent

OpenStudy (lgbasallote):

\[7^{2t-1} = 35\] huh

OpenStudy (lgbasallote):

you take the log

OpenStudy (lgbasallote):

\[\log_7 35 = 2t - 1\]

OpenStudy (lgbasallote):

\[\log_7 35 + 1 = 2t\]

OpenStudy (lgbasallote):

\[\frac{\log_7 35 + 1}{2} = t\]

OpenStudy (anonymous):

using my notes, im supposed to solve it this way 2t log7-log7=log35

OpenStudy (lgbasallote):

oh i see...that's possible...do you want to know where it came from?

OpenStudy (anonymous):

i know how to get to that point but dont know where to go from there

OpenStudy (lgbasallote):

take the log of both sides \[(2t-1)\log 7 = \log 35\] distribute \[2t\log 7 - \log 7 = \log 35\] \[2t \log 7 = \log 35 + \log 7\] \[2t = \frac{\log(35 \times 7)}{\log 7}\] \[t = \frac{\log (35\times 7)}{2\log 7}\]

OpenStudy (lgbasallote):

hah lol

OpenStudy (lgbasallote):

\[t = \frac{\log 245}{2\log 7}\] \[t = \frac 12 \frac{\log 245}{\log 7}\] \[t = \frac 12 \log_7 245\]

OpenStudy (lgbasallote):

\[t = \log_7 \sqrt{245}\]

OpenStudy (lgbasallote):

im not sure if that's same as what i initially wrote though

OpenStudy (anonymous):

the final rounded answer should come to 1.41 i just have no idea how to get to that

OpenStudy (helder_edwin):

if you have a calculator just type if what @lgbasallote wrote first \[ \Large t=\frac{\log 245}{2\log 7}=1.413543... \]